OpenStudy (roadjester):

\[\huge m_pc^2\sqrt{1+{v^2_{m_p}\over c^2}\ \over 1-{v^2_{m_p}\over c^2}}+m_pc^2\over {\sqrt{1+{v^2_M \over 1-{v^2_m \over c^2}}}}\] @LastDayWork think there's a way to simplify that?

OpenStudy (roadjester):

@LastDayWork i didn't know if you knew what gamma was or not so I wrote out the whole thing

OpenStudy (roadjester):

gamma the Lorentz factor

OpenStudy (lastdaywork):

That's how far I could go without using Binomial approximation - \[mc^{2} \frac{ \sqrt{c^2 + v^2} +\sqrt{c^2 - v^2}}{ \sqrt{c^2 + c^2 v^2 - v^2} }\] & I know about gamma, but never remembered its formula..So you did me a favor by substituting it ;)

OpenStudy (lastdaywork):

BTW the v^2 in the denominator is also V(mp)...right ??

OpenStudy (roadjester):

no, there were two m's there were two gammas, the gamma of a particle of mass mp and the gamma of a particle of mass M so the denominator was for the gamma of mass M and therefore velocity M and the other was for the velocity of a mass of mp

OpenStudy (roadjester):

basically it was a relativistic conservation of linear momentum of two particles in in inelastic collision. but one particle was stationary so it went bye bye the mass M was the result and mass mp was the initial moving particle

OpenStudy (roadjester):

there was also conservation of energy; forgot to mention that

OpenStudy (lastdaywork):

In that case, you should directly apply binomial..as I can't see any valid simplification..

OpenStudy (roadjester):

the only question really then is WHERE to begin the binomial...

OpenStudy (roadjester):

\[\huge \sqrt{{(m_pc^2)^2}+{(\gamma_m m_p v_mc)^2}}+m_pc^2 = \huge \sqrt{{(Mc^2)^2}+{(\gamma_M M v_Mc)^2}}\]

OpenStudy (roadjester):

aww, huge is too big

OpenStudy (roadjester):

\[\huge \sqrt{{(Mc^2)^2}+{(\gamma_M M v_Mc)^2}}\]

OpenStudy (roadjester):

where gamma is 1-v^2/c^2 and I'm solving for Mc^2

OpenStudy (lastdaywork):

This question looks too scary to me :( @douglaswinslowcooper is online...maybe he can help here..

OpenStudy (roadjester):

lol

OpenStudy (roadjester):

give me a sec to post the original question then

OpenStudy (anonymous):

@roadjester Where did this algebraic expression come from?

OpenStudy (lastdaywork):

@agent0smith - Can you help here ??

OpenStudy (roadjester):

oh from conservation of energy \[E^2={(mc)}^2+{(pc)}^2\] each particle has it's own energy the letter p is momentum which is \[p=\gamma mv\]

OpenStudy (anonymous):

Okay, so how did you manage to screw it up?

OpenStudy (roadjester):

well that was rude, and I didn't "screw it up" as you so nicely put it

OpenStudy (lastdaywork):

@ybarrap

OpenStudy (anonymous):

There are no egos here. We all all buddies in math.

OpenStudy (roadjester):

this is the physics group bro

OpenStudy (roadjester):

73. The creation and study of new and mery massive elementary particles is an important part of contemporary physics. To create a particle of mass M requires an energy Mc^2. With enough energy, en exotic particle can be created by allowing a fast-moving proton to collide with a similar target particle. Consider a perfectly inelastic collision between two protons: an incident proton with mass m_p, kinetic energy K, and momentum magnitude p joins with an originally stationary target proton to form a single product particle of mass M. Not all the kinetic energy of the incoming proton is available to create the product particle because conservaton of momentum requires that the system as a whole still must have some kintetic energy after the collision. Therefore, only a fraction of the energy of the incident particle is available to create a new particle. (a) Show that the energy available to create a product particle is given by [SEE ATTACHMENT] This result shows that when the kinetic energy K of the incident proton is large compared with its rest energy m_pc^2, then M approaches (2m_pK)^(1/2)/c. Therefore, if the energy of the incoming proton is increased by a factor of 9, the mass you can create increases only by a factor of 3, not by a factor of 9 as would be expected. (b) This problem can be alleviated by using colliding beams as is the case in most modern accelerators. Here the total momentum of a pair of interacting particles can be zero. The center of mass can be at rest after the collision, so, in principle, all the initial kinetic energy can be used for particle creation. Show that [SEE ATTACHMENT] where K is the kinetic energy of each of the two identical colliding particles. Here, if K>>mc^2, we have M directly proportional to K as we would desire.

OpenStudy (anonymous):

You're right. It is physics. I hope you lift.

OpenStudy (anonymous):

Still, at what point did you decide you needed to expand \(\gamma\)?

OpenStudy (lastdaywork):

@roadjester you interpreted @wio wrong..it was meant as a humor.. @ybarrap Can you explain how to simplify the expression in the question...using binomial and stuff..

OpenStudy (ybarrap):

Are you sure this, $$ \huge{ m_pc^2\sqrt{1+{v^2_{m_p}\over c^2}\ \over 1-{v^2_{m_p}\over c^2}}+m_pc^2\over {\sqrt{1+{v^2_M \over 1-{v^2_m \over c^2}}}}}\\ $$ Isn't supposed to be this, $$ \huge{ m_pc^2\sqrt{1+{v^2_{m_p}\over c^2}\ \over 1-{v^2_{m_p}\over c^2}}+m_pc^2\over\sqrt{1+{v^2_{m_p}\over c^2}\ \over 1-{v^2_{m_p}\over c^2}}}\\ $$ ?

OpenStudy (roadjester):

actually there is a mistake in the square roots but it isn't the one you suggested. the 1 in the numerator of the top square root should also be pulled out like the denominator

OpenStudy (roadjester):